OpenStudy (anonymous):

factor each expression. -x^2+3x+40

Directrix (directrix):

-x^2+3x+40 = - ( x² - 3x - 40) = - (x² - 8x + 5x - 40) = - [ x(x - 8) + 5(x - 8) ] = - [(x - 8) (x + 5)] = ? @cutegirl